200g water at 0 degree centigrade mixed with 200g ice at -25 degree centigrade temperature.find the ratio of mass water and ice after mixing.
1
Expert's answer
2020-09-14T10:22:35-0400

Let's suppose that we have ice and water mixture after reaching heat ballance, which is possible only at temperature of 0 celsium. If we have another situation, we will have impossible numbers in equation and will change assumption. Heat equation in this case:

"λm_w=c_im_i\cdot25", where "m_w" is mass of frozen water, and "m_i" is mass of ice. If the water is 0 degree and is freezing, the ice can`t melt. When it reaches 0 degree we will have a system ice-water at 0 degree, which is stable. Thus, "333500\cdot m_w=2050\cdot0.2 \cdot 25 \Rightarrow m_w=\frac{2050\cdot 0.2\cdot 25}{333500}\approx0.031kg".

So, 32 gramms of water will freeze and ratio of water and ice masses is "\frac{200-31}{200+31}=0.73".